# Hierarchical Text Classification  (HTC) vs. eXtreme Multilabel Classification (XML): Two Sides of the Same Medal

### v2

@@ 3. Dataset Transfer XML $\leftrightarrow$ HTC
In this section, we discuss our assumptions and methods for converting datasets between the two worlds of HTC and XML. We describe how the HTC hierarchies are collapsed in order to apply XML methods to these datasets (see Section 3.2), and how the label hierarchy is built for HTC models (see Section 3.3), including the adjustment of model parameters and handling of large datasets that exceed the limits of the encoders.
3.1. Assumptions
The datasets used for our experiments were selected from all the datasets mentioned and used by the different models of HTC and XML for text classification. Thus, it is assumed that the datasets used for our experiment are representative of the domains in which they are typically used. Therefore, we expect our experiment to yield meaningful results.
3.2. Transfer of HTC Datasets for XML Models
@@
This direction is the simpler of the two, therefore the transfer can be described briefly. As the XML methods do not require a predefined label hierarchy, the label hierarchies of the HTC datasets can simply be discarded. After this step, there are two ways to complete the transfer. The first option is to transfer all the labels in the labelset and the second option is to transfer only the labels that appear as leaf nodes in the original hierarchy. These are the most specific labels. Note that depending on the hierarchical dataset, these sets are not necessarily disjoint. Our main experiments focus on the first version. Since HTC models train and predict with internal nodes as well as leaf nodes, we found that this version provides the fairest comparison. Nevertheless, we have included the second version as supplementary material in the appendix.
In addition, CascadeXML and XR-Transformer require model-specific input files that are used during training to create their own hierarchies in the form of an HLT (Hierarchical Label Tree, see Section 2). This creation of an HLT is described in detail in the following Section 3.3. We provide these files in our aforementioned GitHub repository.
3.3. Transfer of XML Datasets for HTC Models
In order for XML datasets to work with the HTC models, a hierarchical label tree as constructed by the XML methods is needed to serve as the hierarchy for the HTC methods. In order to generate the hierarchical label tree, we employ a clustering method similar to CascadeXML. Each label in the set is assigned a numerical value based on the feature aggregation of its positive instances (PIFA) (Kharbanda et al., 2022), which is computed as follows. For a label $l_{i}$ and the set $D_{i}$ of documents labeled with $l_{i}$ in the ground truth, PIFA is defined as:
$\operatorname{PIFA}(l_{i})=\sum_{d\in D_{i}}\frac{\text{TF-IDF}(d)}{\operatorname{length}(d)}.$
@@ 3.3. Transfer of XML Datasets for HTC Models
Subsequently, a balanced $k$-means algorithm is applied recursively on these numeric label representations a total of $n$ times, until a HLT of depth $n$ is created. The difference between our method to transfer XML datasets to HTC models and the application of CascadeXML to the XML datasets is that Equation 3.3 normalizes the TF-IDF vectors by dividing them by the number of words in the documents, while CascadeXML takes the norm of the resulting embedding. This was done to reduce the influence of text length on clustering. This means that labels are clustered together not only when they are relevant to documents with similar features, but also when the labels have similar relevance frequencies. Note, that by generating a hierarchy in this manner, we introduce new labels to the dataset. To allow a meaningful comparison, we use these new intermediate labels in training but remove them during classification. The true labels from the original dataset are now all leaf nodes within the generated hierarchy, even if a label was previously a parent node higher up in the original hierarchy. The branching factors and depths of these trees that were used in our experiments were selected on the basis of CascadeXML. The depth branching factors are summarized in Table 1.
@@
Table 1. Branching factors for each level L1 to L3 (and L4, respectively) of the hierarchical label tree in XML datasets.
@@
Both HBGL and HGCLR use this hierarchy as input to their underlying encoder within the model. In the case of HBGL, this encoder is BERT, while in HGCLR it is Graphormer. Each label (node) in the HLT is reduced to a node embedding and serves as an input token for the encoder. The tree structure is preserved by using a custom attention matrix, whereby nodes can only attend to successive nodes.
The encoders have an inherent limit regarding the maximum number of input tokens. Since each label acts as an input token, all of the XML datasets surpass this limit. Furthermore, the quadratic complexity of the attention mechanism worsens the issue with large input sizes. To address this limitation, we segment the hierarchical label tree into smaller sub-trees, each maintaining the properties of an HLT.
We perform this segmentation using depth-first search (DFS), exclusively traversing downward to gather just the leaf nodes and their corresponding preceding nodes. This method ensures that each leaf node in the original tree remains a leaf node in the new sub-tree. Additionally, the number of labels in each new sub-tree is kept within the token limit of BERT of 512 tokens.
By doing so, we might lose context between certain nodes, as they will not be aware of each other within the attention mechanism of the encoder. However, large parts of the general hierarchical structure should still be preserved, as distant nodes generally interact less with each other, e. g., an article about sports is less likely to contain a lot of keywords related to astrophysics.
@@
4. Experimental Apparatus
In the following, we present datasets, preprocessing, procedure, and performance measures.
4.1. Datasets
In total, we consider six different benchmarking datasets. Three of them are from the world of HTC and three of them are from the world of XML. The HTC datasets are Web of Science - WoS by Kowsari et al. (2017), The New York Times Annotated Corpus - NYT by Shimura et al. (2018) and RCV1-v2 by Lewis et al. (2004). The XML datsets are Wiki10-31K by Zubiaga (2012) and AmazonCat-13K, and Amazon-670K each by McAuley and Leskovec (2013). See Table 2 general datasets statistics. The XML datasets can be obtained from “The Extreme Classification Repository: Multi-Label Datasets & Code” by  Bhatia et al. (2016).
Table 2. Summary statistics of the HTC and XML datasets. Amount of labels and data points within each dataset. Avg($L_{i}$) is the average number of classes per sample.
@@ 4.2. Preprocessing
For the preprocessing of the datasets, we follow the procedures described within the specific models HGCLR, HBGL, XR-Transformer, and CascadeXML. Additionally, when preparing HTC datasets for the XML models and vice versa, we refer back to the dataset transfer described in Section 3. In addition, to apply the HTC models on the XML datasets AmazonCat-13K and Amazon-679k, the number of training examples is set to $30k$ and another $5k$ for validation, while the test set is left unchanged. This is needed to allow for a reasonable training time.
4.3. Procedure
@@
Our first step is the reproduction of published results for all models we use. With the original results confirmed, we move on to convert the three HTC datasets (WoS, NYT, RCV1-V2) by removing the hierarchy and vice versa injecting a hierarchy into the three XML datasets (Wiki10-31K, AmazonCat-13K, Amazon-670K) via clustering. With all preparations done, all four models are run on the respective datasets adjusted to their domain. All results are then averaged over five individual runs.
4.4. Hyperparameter Optimization
For the initial reproduction of the results for each model, we used the hyperparameters provided by the authors, as those have already been tuned. When it comes to our experiment of applying these models to new datasets, we search for optimal hyperparameter values to adapt a model to the new dataset. However, we do so in a constrained manner by focusing on a few parameters that we expect to possibly have a noticeable impact since optimization through all possible hyperparameters is very time and resource-consuming with high complexity. Hyperparameters we consider for optimization are batch size, number of training epochs, and learning rate.
4.5. Performance Measures
The performance of the models is measured differently in the domains of HTC and XML. HTC models typically use the F1-Micro and F1-Macro metrics. While those are based on classification, XML models use the ranking-based metrics Precision@k (P@k) and Propensity Scored Precision@k (PSP@k) (Jain et al., 2016). F1-Micro and F1-Macro are different ways of aggregating the F1 score (Goutte and Gaussier, 2005). F1-Macro gives equal weight to each class regardless of frequency and F1-Micro gives more weight to frequent classes. F1-Marco is calculated by averaging the F1 score over every class and F1-Micro is calculated by treating every instance equally and computing the harmonic mean of the global (micro) precision and recall across all classes. P@k measures how well a model accurately identifies relevant elements within its top-$k$ predictions. It quantifies how many top-$k$ predictions are also marked as relevant in the ground truth. P@k is formally defined as
$\operatorname{P@k}=\frac{1}{k}\sum_{i=1}^{k}\text{Relevance}(i)$
where Relevance(i) is an indicator function that is $1$ if the $i$-th item is relevant and 00 otherwise. PSP@k does the same, but has a special emphasis on tail labels and rewards their accurate prediction, as they are often challenging to predict (Jain et al., 2016). Tail labels refer to those labels that occur infrequently across the dataset. These labels are part of the long tail distribution, where a few labels are very common (head labels), and many labels are rare (tail labels) (Yue et al., 2024). This metric is particularly relevant in the world of XML because the labelsets are very large, which means that tail labels occur often. In order to enable a meaningful comparison of the performance of models from the HTC and XML domains, it is necessary to define a common evaluation metric. In this context, we decided to adopt the P@k metric from the XML domain and implement it for the HTC models. Furthermore, we use the R-Precision metric, which is an adaptation of the P@k metric, that does not set a fixed $k$ but rather sets $k$ to the number of relevant labels in the ground truth for every data point that is being evaluated (Manning et al., 2009). Formally, R-Precision is defined as
$\operatorname{R\text{-}Precision}=\frac{1}{r}\sum_{i=1}^{r}\text{Relevance}(i)$
where $r$ is the number of relevant labels in the ground truth for the data point that is evaluated.
It is possible to implement these ranking-based metrics for the classification-based HTC models because they work by assigning each label a probability of being relevant to the current query and then classifying by checking if this probability exceeds a certain threshold. This means that they internally also employ a ranking, which can be utilized to calculate P@k scores and the R-Precision. Another possible way of establishing a common metric would have been to adopt the ranking-based XML models to make classification decisions, which would allow the calculation of F1 scores. To do this, one would need to take the relevance ranking produced by the XML models and set a fixed cut-off point above which labels are classified as relevant. However, setting this cut-off point is not trivial and would need to be optimized separately for each dataset. It should also be noted that it is not useful to use a dynamic cut-off point. For example, taking the approach of the R-Precision metric and setting the cut-off point to the number of relevant labels in the ground truth for each query would mean that precision and recall are equivalent, which would make the calculation of F1 scores obsolete (Manning et al., 2009). Furthermore, we excluded the PSP@k metric, due to the considerably smaller size of the HTC labelsets in comparison to the XML labelsets. This makes the problem with the tail labels less relevant for the HTC methods.
The training time of the different models is not taken into consideration in this paper due to the lack of information provided in the original papers and because we had to vary the hardware configuration during experimentation to accommodate specific model-dataset combinations. However, it can be said that XML models generally train substantially faster than HTC models.
5. Results
We report the performance of HTC and XML models on datasets originating from the HTC world. Subsequently, we report the results on datasets from the XML world. The results are the mean of five individual runs for each model-dataset combination.
5.1. HTC and XML methods on HTC datasets
@@
Table 3. Results for the HTC datasets, namely New York Times Annotated Corpus (top), Web of Science (middle), Reuters Corpus Volume 1 (bottom). HBGL** indicates that the semantic hierarchy, as provided by the datasets in HTC, has been replaced by a synthetic hierarchical label tree as it is computed in XML.
@@
The results on the WoS dataset (see Table 3, middle) show a different picture. HGCLR performs best across all metrics, outperforming HBGL by $0.42$ points, CascadeXML by $3.06$ points, and XR-Transformer by $13.46$ points in R-Precision, respetively. In contrast to the WoS and RCV1 datasets, HBGL has the second-highest R-precision, beating CascadeXML by $2.64$ points. XR-Transformer is well behind all other methods. Note the special nature of the WoS dataset. Each data point has exactly two associated labels, so the precision values for P@2 and R-precision are the same (see Table 2 for details).
The results for the RCV1-V2 dataset (see Table 3, bottom) are similar to those for the NYT dataset. HGCLR performs best on P@2, P@3, P@5, and R-precision with a small margin over CascadeXML. Regarding P@1, CascadeXML has the upper hand with a difference of $0.53$ points to HGCLR. For the other metrics, the improvements of HGCLR over CascadeXML vary from $0.51$ (P@2) to $1.48$ (R-Precision). The differences between CascadeXML and XR-Transformer are rather small for all metrics ($<1.0$ point) and HBGL is outperformed by both methods. Again, the results of HGCLR and CascadeXML are close together, followed by XR-Transformer and HBGL.
5.2. HTC and XML methods on XML datasets
@@
Table 4. Results for the XML datasets, namely Wiki10-31K (top), AmazonCat-13K (middle), and Amazon-670K (bottom). For methods marked with † no results could be obtained due to resource limitations.
@@
Results on the AmazonCat-13K dataset (results are shown in Table 4, middle) show a slightly different picture in the comparison of the XML models. CascadeXML outperforms both XR-Transformer and HBGL, reaching an R-Precision of $83.13$. Although the improvement upon XR-Transformer is rather small with $1.11$ points, again a sizeable gap to HBGL is observed with a difference of $28.46$ points in R-Precision. Also, the overall performance for this dataset is high compared to other datasets, reaching values in R-Precision up to $83.13$.

### v3

@@ 3.1 Bidirectional Transfer of Datasets
To answer our main research question of a bidirectional transfer of methods between HTC and XML, we need to convert the datasets to make the respective other methods applicable. In the following, we describe how the HTC hierarchies are collapsed in order to apply XML methods to these datasets. The other direction of constructing a label hierarchy for HTC models to work with XML datasets is described in Section 3.3, including adjustment of model parameters and handling of large datasets exceeding the limits of the models.
@@
The XML methods do not require a pre-defined label hierarchy so that the pre-defined hierarchy of the HTC can simply be discarded. After this step, however, there are two ways to complete the transfer. The first option is to transfer all the labels in the label set, and the second option is to transfer only the labels that appear as leaf nodes in the original hierarchy. These are the most specific labels. Note that depending on the hierarchical dataset, these sets are not necessarily disjoint. Our main experiments focus on the first version. Since HTC models train and predict with internal nodes as well as leaf nodes, we find that this version is most suitable to facilitate a fair comparison (no changes to the ground truth of assigned labels). In addition, CascadeXML and XR-Transformer need to construct a hierarchical label tree (see Section 2) for these datasets in a preprocessing step. How exactly this is done depends on the specific models. The construction of such hierarchical label trees is described in detail in Section 3.3.
3.2 Harmonizing Performance Metrics
The performance of the models is measured differently in the domains of HTC and XML. HTC models typically use the classification metrics F1-Micro and F1-Macro [10]. F1-Micro and F1-Macro are different ways of aggregating the F1 score [10]. F1-Macro gives equal weight to each class regardless of frequency, and F1-Micro gives more weight to frequent classes. F1-Marco is calculated by averaging the F1 score over every class, and F1-Micro is calculated by treating every instance equally and computing the harmonic mean of the global (micro) precision and recall across all classes.
In contrast, XML models use the ranking-based metrics Precision@k (P@k) and Propensity Scored Precision@k (PSP@k) [12]. P@k measures how well a model accurately identifies relevant elements within its top-$k$ predictions. It quantifies how many top-$k$ predictions are also marked as relevant in the ground truth. P@k is formally defined as $\operatorname{P@k}=\frac{1}{k}\sum_{i=1}^{k}\text{Relevance}(i)$ where Relevance(i) is an indicator function that is $1$ if the $i$-th item is relevant and 0 otherwise. PSP@k does the same, but has a special emphasis on tail labels and rewards their accurate prediction, as they are often challenging to predict [12]. Tail labels refer to those labels that occur infrequently across the dataset [40]. This metric is particularly relevant in research on XML because the label sets are very large, which means that tail labels occur often.
For a meaningful comparison of the performance of models across the HTC and XML domains, it is necessary to define a common evaluation metric. An option is the R-Precision metric, which is an adaptation of the P@k metric that does not set a fixed $k$ but rather sets $k$ to the number of relevant labels in the ground truth for every data point that is being evaluated [21]. It is defined as $\operatorname{R\text{-}Precision}=\frac{1}{r}\sum_{i=1}^{r}\text{Relevance}(i)$ where $r$ is the number of relevant labels in the ground truth for the data point that is evaluated. The metric is appealing as one does not need to consider multiple values of $k$ as in $P@k$. Calculating these ranking-based metrics for the classification-based HTC models is usually possible because the methods learn probabilities for each label. This means that they internally also employ a ranking (e.g., it can be extracted from the logits), which can be utilized to calculate P@k scores and the R-Precision.
Another possible way of establishing a common metric would have been to modify the ranking-based XML models to make hard classification decisions (whether one specific class will be assigned or not), which would allow the calculation of classification metrics such as the F1 score. To do this, one would need to take the relevance ranking produced by the XML models and determine the cut-off point above which labels are classified as relevant. However, setting this cut-off point is not trivial and would need to be optimized separately for each dataset. In fact, this conversion of ranking to hard decisions is a research endeavor of its own, e.g., see Tang et al. [29], but usually not considered in XML papers. It should also be noted that it is not feasible to use a dynamic cut-off point based on the ground truth because this would not be available when the model is deployed. Even if one would take the approach, for example, of using the R-Precision metric and setting the cut-off point to the number of relevant labels in the ground truth for each query, that would entail that precision and recall would be mathematically equivalent, which would make the calculation of F1 scores obsolete [21].
For these reasons, we decided to consistently employ the ranking-based metrics $\operatorname{P@k}$ and $\operatorname{R-Precision}$ to facilitate our comparison. We excluded the PSP@k metric, due to the considerably smaller size of the label sets in HTC in comparison to the label sets in XML. This makes the problem of tail labels less relevant for the HTC methods.
3.3 Transfer of XML Datasets for HTC Models
@@
In order for HTC methods to work with XML datasets, a hierarchical label tree as constructed by the XML methods is needed to serve as the hierarchy for the HTC methods. In order to generate the hierarchical label tree, we employ a clustering method similar to CascadeXML. Each label in the set is assigned a numerical value based on the feature aggregation of its positive instances (PIFA) [16], which is computed as follows. For a label $l_{i}$ and the set $D_{i}$ of documents labeled with $l_{i}$ in the ground truth, PIFA is defined as $\operatorname{PIFA}(l_{i})=\sum_{d\in D_{i}}\frac{\text{TF-IDF}(d)}{\operatorname{length}(d)}$. Subsequently, a balanced $k$-means algorithm is applied recursively on these numeric label representations a total of $n$ times until a hierarchical label tree of depth $n$ is created. The difference between our method to transfer XML datasets to HTC models and the application of CascadeXML to the XML datasets is that we normalize the TF-IDF vectors by dividing them by the number of words in the documents, while CascadeXML takes the norm of the resulting embedding. This was done to reduce the influence of text length on clustering. This means that labels are clustered together not only when they are relevant to documents with similar features but also when the labels have similar relevance frequencies.
Note that by generating a hierarchy in this manner, we introduce new labels to the dataset. To allow a meaningful comparison, we use these new intermediate labels in training but remove them during classification. The true labels from the original dataset are now all leaf nodes within the generated hierarchy, even if a label was previously a parent node higher up in the original hierarchy. The branching factors and depths of these trees that were used in our experiments were selected on the basis of CascadeXML. The depth branching factors are summarized in Table 1.
@@
Table 1: Branching factors for each level L1 to L3 (and L4, respectively) of the hierarchical label tree in XML datasets.
@@
Both HBGL and HGCLR use this hierarchy as input to their underlying encoder within the model. In the case of HBGL, this encoder is BERT, while in HGCLR it is Graphormer. Each label (node) in the hierarchical label tree is reduced to a node embedding and serves as an input token for the encoder. The tree structure is preserved by using a custom attention matrix, whereby nodes can only attend to successive nodes.
The encoders have an inherent limit regarding the maximum number of input tokens. Since each label acts as an input token, all of the XML datasets surpass this limit. Furthermore, the quadratic complexity of the attention mechanism worsens the issue with large input sizes. To address this limitation, we segment the hierarchical label tree into smaller sub-trees, each maintaining the properties of a hierarchical label tree.
We perform this segmentation using depth-first search (DFS), exclusively traversing downward to gather just the leaf nodes and their corresponding preceding nodes. This method ensures that each leaf node in the original tree remains a leaf node in the new sub-tree. Additionally, the number of labels in each new sub-tree is kept within the token limit of BERT of 512 tokens. By doing so, we might lose context between certain nodes, as they will not be aware of each other within the attention mechanism of the encoder. However, large parts of the general hierarchical structure should still be preserved, as distant nodes generally interact less with each other, e. g., an article about sports is less likely to contain a lot of keywords related to astrophysics.
@@
To apply the HTC models on the XML datasets AmazonCat-13K and Amazon-679k, the number of training examples is set to $30k$ and another $5k$ for validation, while the test set is left unchanged. This is needed to allow for a reasonable training time.
4 Experimental Apparatus
Datasets
We consider six benchmark datasets, three HTC and three XML, see Table 2. The HTC datasets are Web of Science (WoS) [17], The New York Times Annotated Corpus (NYT) [26], and RCV1-v2 [18]. Commonly used XML datasets are Wiki10-31K [46], as well as AmazonCat-13K and Amazon-670K [22]. All datasets are available publicly or after signing a license; NYT, while no longer licensable, is used under an existing valid license. We use NYT because it is frequently applied in HTC research.
Table 2: Summary statistics of the HTC and XML datasets. Amount of labels and data points within each dataset. Avg($L_{i}$) is the average number of classes.
@@ 4 Experimental Apparatus
For preprocessing, we follow the procedures described within the specific models HGCLR, HBGL, XR-Transformer, and CascadeXML. When preparing HTC datasets for the XML models and vice versa, we refer back to the dataset transfer described in Section 3.
Procedure
@@
Our first step is the reproduction of published results for all models we use. Second, we convert the HTC datasets (WoS, NYT, RCV1-V2) by removing the hierarchy and vice versa, injecting a hierarchy into the XML datasets (Wiki10-31K, AmazonCat-13K, Amazon-670K) via clustering. With all preparations done, all four models are run on the respective datasets adjusted to their domain. Results are averaged over five runs.
Hyperparameter Optimization
For the reproduction of the results for each model, we used the hyperparameters provided by the authors, as those have already been tuned. When transferring methods to the datasets of the respective other domains, we chose the hyperparameters of the most similar datasets. For the XML methods on HTC datasets, we use those hyperparameters that each respective paper reports for the smallest XML dataset: Wiki10-31K.
5 Results
We report the performance of HTC and XML models on datasets from the HTC stream of research. The results for the other direction, i. e., applying HTC and XML models to XML datasets can be found in Section 5.2.
5.1 HTC and XML methods on HTC datasets
@@
Table 3: Results for the HTC datasets, namely New York Times Annotated Corpus (top), Web of Science (middle), Reuters Corpus Volume 1 (bottom). HBGL** indicates that the semantic hierarchy, as provided by the datasets in HTC, has been replaced by a synthetic hierarchical label tree as it is computed in XML.
@@
The results on the WoS dataset (see Table 3, middle) show a different picture. HGCLR performs best across all metrics, outperforming HBGL by $0.42$ points, CascadeXML by $3.06$ points, and XR-Transformer by $13.46$ points in R-Precision, respectively. In contrast to the WoS and RCV1 datasets, HBGL has the second-highest R-precision, beating CascadeXML by $2.64$ points. XR-Transformer is well behind all other methods. Note that for WoS, each data point has exactly two labels (see Table 2), so the P@2 equals the R-precision .
The results for the RCV1-V2 dataset (see Table 3, bottom) are similar to those for the NYT dataset. HGCLR performs best on P@2, P@3, P@5, and R-precision with a small margin over CascadeXML. Regarding P@1, CascadeXML has the upper hand with a difference of $0.53$ points to HGCLR. For the other metrics, the improvements of HGCLR over CascadeXML vary from $0.51$ (P@2) to $1.48$ (R-Precision). The differences between CascadeXML and XR-Transformer are small for all metrics (<1.0<1.0 point) and HBGL is outperformed by both methods. Again, the results of HGCLR and CascadeXML are close, followed by XR-Transformer and HBGL.
5.2 HTC and XML methods on XML datasets
@@
Table 4: Results for the XML datasets, namely Wiki10-31K (top), AmazonCat-13K (middle), and Amazon-670K (bottom). For methods marked with † no results could be obtained due to resource limitations.
@@
Results on the AmazonCat-13K dataset (see Table 4, middle) show a slightly different picture in the comparison of the XML models. CascadeXML outperforms both XR-Transformer and HBGL, reaching an R-Precision of $83.13$. Although the improvement upon XR-Transformer is rather small with $1.11$ points, again, a sizeable gap to HBGL is observed with a difference of $28.46$ points in R-Precision. Also, the overall performance for this dataset is high compared to other datasets, reaching values in R-Precision up to $83.13$.